# Fearless Stochasticity in Expectation Propagation

### v1

@@ 4.2 Cosmic radiation model
A hierarchical Bayesian model was used by Vehtari et al. [44] to capture the nonlinear relationship between diffuse galactic far ultraviolet radiation and 100-µm infrared emission in various sectors of the observable universe, using data obtained from the Galaxy Evolution Explorer telescope. In this model each $w_{i}\in\mathbb{R}^{9}$ parameterised a nonlinear regression model using data obtained from one of $m$ sections of the observable universe. The $m$ regression problems were related through hyperparameters $z\in\mathbb{R}^{18}$, which parameterised the section-level coefficient prior densities $p(w_{i}\mid z)$ for $i=1,\ldots,m$. $p_{0}\in\mathcal{F}$ was MVN. The specifics of the nonlinear regression model are quite involved and we refer the reader to Vehtari et al. [44] for details.
Vehtari et al. [44] used EP to perform inference in this model. We were unable to obtain the dataset, and so we generated synthetic data using hyperparameters parameters that were tuned by hand to try and match the qualitative properties of the original data; see Appendix K for examples. We used a reduced number of $m=36$ sites and $n=200$ observations per site to allow us to perform a comprehensive hyperparameter search. Results are shown in Figure 2c.
4.3 Neural response model
@@
A common task in neuroscience is to model the firing rates of neurons under various conditions. We performed inference in a hierarchical Bayesian neural response model, using recordings of V1 complex cells in an anesthesised adult cat [3]. In this dataset, 10 neurons in a specific area of cat V1 were simultaneously recorded under the presentation of 18 different visual stimuli, each repeated 8 times, for a total of 144 trials. We modelled the observed spike counts of the 10 neurons in each trial as Poisson, with latent (log) intensities being MVN, with mean and covariance (collectively $z\in\mathbb{R}^{65}$) drawn from a NIW prior $p_{0}\in\mathcal{F}$. Inference of $z$ amounted to inferring the means, and inter-neuron correlations, of latent log firing rates across stimuli. We grouped the data into $m=8$ batches of $n=18$ trials each, so that the latent log firing rates for all trials in batch $i$ were jointly captured by $w_{i}\in\mathbb{R}^{180}$. Results are shown in Figure 2d.
@@ 5 Conclusion
Limitations
We largely assumed that the drawing of MC samples is the dominant cost. In practice, other costs become relevant, which may shift the balance in favour of using more samples to estimate updates. We discuss this in Appendix G, along with strategies for reducing computational overheads. We note, however, that wall-clock time results for our experiments (in Appendix J) are in line with the results of Section 4. When EP and its variants—including those introduced here—are combined with MC methods, the underlying samplers typically have hyperparameters of their own, which are often adapted using so-called warmup phases. While EP-$\eta$ and EP-$\mu$ significantly reduce the complexity of tuning hyperparameters specific to EP, they do not help with those of the underlying samplers; this means that the complexity of hyperparameter tuning is still greater than that of direct MC methods. We used comprehensive hyperparameter searches in our experiments in order to perform a meaningful comparison with baseline methods, necessarily limiting the scale of problems we could tackle.
Future work
It would be worthwhile considering how our algorithms can be efficiently implemented for modern parallel compute hardware; e.g., practical performance could be improved by using sampling kernels that are themselves designed for efficient parallel execution [23]. Performance may also be improved in distributed settings by using asynchronous extensions to minimise communication overheads [14, 47]. The need to tune sampling hyperparameters could be obviated by using information from $p$ to set them directly; e.g., when $p\in\mathcal{F}$ is MVN, its precision matrix could be used directly as the mass matrix hyperparameter in Hamiltonian Monte Carlo schemes. We would also like to apply our methods to large-scale problems. We leave investigation of these ideas to future work.
@@ Appendix A Exponential families of distributions
In this appendix, we give a very brief overview of exponential families of distributions. This is a highly condensed summary of relevant material from Wainwright and Jordan [45], and we refer the reader to the original work for a far more detailed treatment.
The exponential family of distributions $\mathcal{F}$, defined by the ($d$-dimensional, vector-valued) sufficient statistic function $s(.)$ and base measure $\nu(.)$, has density function
$p(z)$ $=\exp\bm{(}\theta^{\top}s(z)-A(\theta)\bm{)},$ (16)
taken with respect to $\nu(.)$, where $A(\theta)=\log\int\exp\bm{(}\theta^{\top}s(z)\bm{)}\mathrm{d}\nu(z)$ is known as the log partition function. $\theta\in\Omega$ are known as the natural parameters, and are used to index a specific member of $\mathcal{F}$. $\Omega$ is the set of normalisable natural parameters of $\mathcal{F}$, given by
$\Omega=\left\{\theta\in\mathbb{R}^{d}\ \Bigg{\lvert}\ \int\exp\bm{(}\theta^{\top}s(z)\bm{)}\mathrm{d}\nu(z)<\infty\right\},$ (17)
and is known as the natural domain of $\mathcal{F}$. $\Omega$ is a convex set, and when it is open, the family $\mathcal{F}$ is said to be regular; we shall only consider regular families in this work. When the components of $s(.)$ are linearly independent, $\mathcal{F}$ is said to be minimal. For minimal families, each distribution in the family is associated with a unique natural parameter vector $\theta$. An exponential family that is not minimal is said to be overcomplete, in which case, each distribution is associated with an entire affine subspace of $\Omega$.
The expected sufficient statistics of a distribution with density $p(.)$ with respect to base measure $\nu(.)$, are given by $\mathbb{E}_{p(z)}[s(z)]$. The mean domain, $\mathcal{M}$, is defined as the set of expected sufficient statistics that can be attained by any density $p(.)$ with respect to base measure $\nu(.)$; that is,
@@
All vectors in the interior of $\mathcal{M}$, $\mathcal{M}^{\circ}$, are realisable by a member of $\mathcal{F}$, and so $\mu\in\mathcal{M}^{\circ}$ provides an alternative parameterisation of $\mathcal{F}$, in which $\mu$ are known as the mean parameters.
@@
$A^{*}(\mu)=\max_{\theta\in\Theta}\theta^{\top}\mu-A(\theta),$ (19)
and on $\mathcal{M}^{\circ}$, $A^{*}(\mu)$ is equal to the negative entropy of the member of $\mathcal{F}$ with mean parameter $\mu$, hence $A^{*}(.)$ is also referred to as the negative entropy function.
The mean parameters of the member of $\mathcal{F}$ with natural parameter $\theta$ can be obtained by $\mu=\nabla A(\theta)$. For minimal families, this mapping is one-to-one, and the reverse map is given by $\theta=\nabla A^{*}(\theta)$. For this reason, $\nabla A(.)$ and $\nabla A^{*}(.)$ are sometimes referred to as the forward mapping and backward mapping respectively. In this work, we say that a family $\mathcal{F}$ is tractable if both the forward and backward mappings can be evaluated efficiently.
The Fisher information matrix (FIM) of an exponential family distribution, with respect to natural parameters $\theta$, is given by, $F(\theta)=\nabla^{2}A(\theta)$. Furthermore, in minimal families, the FIM with respect to mean parameters $\mu$ is given by $F(\mu)=\nabla^{2}A^{*}(\mu)$. Using the forward and backward mappings, we also have
$F(\theta)$ $=\partialderivative{\mu}{\theta}$ (20)
$F(\mu)$ $=\partialderivative{\theta}{\mu},$ (21)
@@
$F(\theta)^{-1}$ $=\partialderivative{\theta}{\mu}$ (22)
$F(\mu)^{-1}$ $=\partialderivative{\mu}{\theta}.$ (23)
Appendix B EP inner update derivation
@@
$\nabla_{\lambda_{i}}L(\eta_{1},\ldots,\eta_{m},\lambda_{1},\ldots,\lambda_{m})$ $=\nabla A\left(\eta_{0}+\sum_{j}\lambda_{j}\right)-\nabla_{\lambda_{i}}A_{i}\bm{(}(\eta_{i}-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}.$ (24)
@@
$\nabla A\left(\eta_{0}+\lambda_{i}^{\prime}+\sum_{j\neq i}\lambda_{j}\right)$ $=\nabla_{\lambda_{i}}A_{i}\bm{(}(\eta_{i}-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}$
@@
Appendix C Natural gradient view of EP
Let $\tilde{p}_{i}^{(t)}$ be the member of $\mathcal{F}$ with natural parameter $\eta_{i}^{(t)}(\lambda_{i})=\theta_{0}+\sum_{j}\lambda_{j}^{(t)}+\alpha^{-1}(\lambda_{i}-\lambda_{i}^{(t)})$. Proposition 1 states that that the moment-matching updates of EP can be viewed as performing NGD in $L$ with respect to the mean parameters, $\mu_{i}$, of $\tilde{p}_{i}^{(t)}$. We restate the proposition below, and give a proof. Note that the right hand side of (7) maps the EP update (4) onto mean parameters of $\tilde{p}_{i}^{(t)}$. The left hand side of (7) is simply a NGD update in $\mu_{i}$.

### v2

@@ 4 Evaluation
A hierarchical Bayesian model was used by Vehtari et al. [48] to capture the nonlinear relationship between diffuse galactic far ultraviolet radiation and 100-µm infrared emission in various sectors of the observable universe, using data obtained from the Galaxy Evolution Explorer telescope. In this model each $w_{i}\in\mathbb{R}^{9}$ parameterised a nonlinear regression model using data obtained from one of $m$ sections of the observable universe. The $m$ regression problems were related through hyperparameters $z\in\mathbb{R}^{18}$, which parameterised the section-level parameter densities. The prior, $p_{0}\in\mathcal{F}$, was MVN. The specifics of the nonlinear regression model are quite involved and we refer the reader to Vehtari et al. [48] for details. We were unable to obtain the dataset, and so we generated synthetic data using parameters that were tuned by hand to try and match the qualitative properties of the original data – see Appendix N for examples. We used a reduced number of $m=36$ sites and $n=200$ observations per site to allow us to perform a comprehensive hyperparameter search. The lower-left panel of Figure 2 shows once again that EP-$\eta$ and EP-$\mu$ consistently reach a given level of accuracy significantly faster than EP and SNEP.
Neural response model
@@
A common task in neuroscience is to model the firing rates of neurons under various conditions. We performed inference in a hierarchical Bayesian neural response model, using recordings of V1 complex cells in an anesthesised adult cat [4]. In this dataset, 10 neurons in a specific area of cat V1 were simultaneously recorded under the presentation of 18 different visual stimuli, each repeated 8 times, for a total of 144 trials. We modelled the observed spike counts of the 10 neurons in each trial as Poisson, with latent (log) intensities being MVN, with mean and covariance (collectively $z\in\mathbb{R}^{65}$) drawn from a NIW prior $p_{0}\in\mathcal{F}$. Inference of $z$ amounted to inferring the means, and inter-neuron covariances, of latent log firing rates across stimuli. We grouped the data into $m=8$ batches of $n=18$ trials each, so that the latent log firing rates for all trials in batch $i$ were jointly captured by $w_{i}\in\mathbb{R}^{180}$. The lower-right panel of Figure 2 shows that EP-$\eta$ and EP-$\mu$ either match or beat the speed of EP for reaching any given level of accuracy. For high levels of accuracy the performances of the methods appear to converge with one another, but note that the asymptote is not exactly zero because the optimal solution was estimated by running EP with a large (but finite) number of samples.
5 Limitations
In this work we have largely assumed that the drawing of MC samples is the dominant cost. In practice, other costs become relevant, which may shift the balance in favour of using more samples to estimate updates. We discuss this in Appendix I, along with strategies for reducing computational overheads. We note, however, that wall-clock time results for our evaluation experiments (in Appendix L) are in line with the results of Section 4, indicating that the drawing of MC samples was indeed the dominant cost.
When EP and its variants—including those introduced here—are combined with MC methods, the underlying samplers typically have hyperparameters of their own, which are often adapted using so-called warmup phases. While EP-$\eta$ and EP-$\mu$ significantly reduce the complexity of tuning hyperparameters specific to EP, they do not help with those of the underlying samplers. This means that the complexity of hyperparameter tuning (for all EP variants) is still greater than that for direct MC methods. We used comprehensive hyperparameter searches in our experiments in order to perform a meaningful comparison with baseline methods, necessarily limiting the scale of problems we could tackle.
Figure 3: Comparison of EP-$\eta$ with conjugate-computation variational inference (CVI) on a hierarchical logistic regression model. The two leftmost plots show forward (solid) and reverse (dashed) KL divergences between the approximation of each method and a MVN distribution estimated directly from MCMC samples. The left panel shows a comparison with respect to wall-clock time, when NUTS is used as the underlying sampling kernel for EP-$\eta$. The left-middle panel shows a similar comparison, but with respect to the number of samples drawn, and using an “oracle” sampling kernel for EP-$\eta$. Hyperparameters were tuned for each method and shaded regions show the range of trajectories across 5 random seeds. The right and right-middle panels show pairwise marginals of the various MVN approximations overlaid on contours of the true posterior. Coloured dots and ellipses correspond to means and 2-standard-deviation contours, respectively. See Section 5 for discussion of these results, and Appendix M for further details.
Our focus has been on developing improved methods for performing EP when the updates must be estimated with noise, with the motivation for this kind of setting having been discussed by prior work [52, 17, 48]. EP and its variants are agnostic to the choice of sampling kernel, and as such we have not discussed issues relating to the performance of the underlying samplers. In the left panel of Figure 3 however, we highlight the relative inefficiency of using EP-$\eta$ with NUTS as the underlying sampler when compared to conjugate-computation variational inference (CVI), an efficient variational inference method [31]. While EP-$\eta$ is able to obtain much more faithful posterior approximations than CVI, it is significantly slower when measured in wall-clock time, despite the gains made over its predecessors. However, this is due to the relative cost of drawing independent samples with NUTS, as the sample efficiency of EP-$\eta$ is roughly equivalent to that of CVI when an (illustrative) “oracle” sampling kernel is used, as demonstrated in the left-middle panel of Figure 3. Further discussion and details of the experiments behind Figure 3 can be found in Appendix M.
6 Future work
In Section 5, we demonstrated that the efficiency of EP variants is constrained by that of of the underlying samplers. There are several ways in which their performance could be improved by using knowledge of $p$ to guide sampling from $p_{i}$. One such improvement would be to use $p$ to set MCMC hyperparameters directly, obviating the need to adapt them during warm-up phases, e.g. when $p\in\mathcal{F}$ is MVN, its precision matrix could be used directly as the mass matrix in Hamiltonian Monte Carlo schemes. Another potential improvement is that samples from $p$ could be used to re-initialise MCMC chains at each iteration, providing approximate independence between updates for a small additional cost. It would also be worthwhile considering how our methods can be efficiently implemented on modern compute hardware to better take advantage of the inherently parallel nature of EP, e.g. practical performance could be improved by using sampling kernels that are themselves designed for efficient parallel execution [26]. Performance may be further improved in distributed settings by using asynchronous extensions of EP in order to minimise communication-related overheads and delays [17, 52]. We leave investigation of these ideas to future work.
@@ 7 Conclusion
8 Acknowledgements
We would like to thank the anonymous reviewers for providing valuable feedback, and Jihao Andreas Lin for sharing his plotting expertise. Jonathan So is supported by the University of Cambridge Harding Distinguished Postgraduate Scholars Programme. Richard E. Turner is supported by Google, Amazon, ARM, Improbable, EPSRC grant EP/T005386/1, and the EPSRC Probabilistic AI Hub (ProbAI, EP/Y028783/1).
@@ Appendix A Exponential families of distributions
In this appendix, we give a very brief overview of exponential families of distributions. This is largely a condensed summary of relevant material from Wainwright and Jordan [49], and we refer the reader to the original work for a far more detailed treatment.
The exponential family of distributions $\mathcal{F}$, defined by the $d$-dimensional, vector-valued sufficient statistic function $s(.)$, and base measure $\nu(.)$, has density function
$p(z)$ $=\exp\bm{(}\eta^{\top}s(z)-A(\eta)\bm{)},$ (16)
taken with respect to $\nu(.)$, where $A(\eta)=\log\int\exp\bm{(}\eta^{\top}s(z)\bm{)}\mathrm{d}\nu(z)$ is known as the log-partition function. $\eta\in\Omega$ are known as the natural parameters, and are used to index a specific member of $\mathcal{F}$. $\Omega$ is the set of normalisable natural parameters of $\mathcal{F}$, given by
$\Omega=\left\{\eta\in\mathbb{R}^{d}\ \Bigg{\lvert}\ \int\exp\bm{(}\eta^{\top}s(z)\bm{)}\mathrm{d}\nu(z)<\infty\right\},$ (17)
and is known as the natural domain of $\mathcal{F}$. $\Omega$ is a convex set, and when it is open, the family $\mathcal{F}$ is said to be regular – we shall only consider regular families in this work. When the components of $s(.)$ are linearly independent, $\mathcal{F}$ is said to be minimal. For minimal families, each distribution in the family is associated with a unique natural parameter vector $\eta$. An exponential family that is not minimal is said to be overcomplete, in which case, each distribution is associated with an entire affine subspace of $\Omega$.
The expected sufficient statistics of a distribution with density $p(.)$ with respect to base measure $\nu(.)$, are given by $\mathbb{E}_{p(z)}[s(z)]$. Let $\mathcal{M}$ be defined as the set of expected sufficient statistics that can be attained by any density $p(.)$ with respect to base measure $\nu(.)$ – that is,
@@
All vectors in the interior of $\mathcal{M}$, $\mathcal{M}^{\circ}$, are realisable by a member of $\mathcal{F}$, and so $\mu\in\mathcal{M}^{\circ}$ provides an alternative parameterisation of $\mathcal{F}$, in which $\mu$ are known as the mean parameters, and $\mathcal{M}^{\circ}$ is known as the mean domain,.
@@
$A^{*}(\mu)=\sup_{\eta\in\Omega}\eta^{\top}\mu-A(\eta),$ (19)
and on $\mathcal{M}^{\circ}$, $A^{*}(\mu)$ is equal to the negative entropy of the member of $\mathcal{F}$ with mean parameter $\mu$.
The mean parameters of the member of $\mathcal{F}$ with natural parameter $\eta$ can be obtained by $\mu=\nabla A(\eta)$. For minimal families, this mapping is one-to-one, and the reverse map is given by $\eta=\nabla A^{*}(\mu)$. For this reason, $\nabla A(.)$ and $\nabla A^{*}(.)$ are sometimes referred to as the forward mapping and backward mapping respectively. In this work, we say that a family $\mathcal{F}$ is tractable if both the forward and backward mappings can be evaluated efficiently.
The Fisher information matrix (FIM) of an exponential family distribution, with respect to natural parameters $\eta$, is given by, $F(\eta)=\nabla^{2}A(\eta)$. Furthermore, in minimal families, the FIM with respect to mean parameters $\mu$ is given by $F(\mu)=\nabla^{2}A^{*}(\mu)$. Using the forward and backward mappings, we also have
$F(\eta)$ $=\partialderivative{\mu}{\eta}$ (20)
$F(\mu)$ $=\partialderivative{\eta}{\mu},$ (21)
@@
$F(\eta)^{-1}$ $=\partialderivative{\eta}{\mu}$ (22)
$F(\mu)^{-1}$ $=\partialderivative{\mu}{\eta}.$ (23)
Appendix B Conventional presentation of EP updates
In this appendix we give a more conventional presentation of the EP updates, and show equivalence with the atypical presentation of Section 2. We will assume here that $\beta_{i}=1\;\forall\;i$, but similar reasoning can be used to show equivalence in the power EP case, where $\beta_{i}\neq 1$. Note that the update of power EP is often presented as having scale proportional to $\beta_{i}$, with a separate damping parameter. We subsume both into $\alpha$, with $\alpha=1$ corresponding to the default damping suggested by Minka [34].
Given a target distribution with the form
$p^{*}(z)$ $\propto\exp\big(\eta_{0}^{\top}s(z)\big{missing})\prod_{i=1}^{m}\exp\big(\ell_{i}(z)\big{missing}),$ (24)
EP attempts to find an approximation $p\in\mathcal{F}$ such that
$p(z)$ $\propto\exp\big(\eta_{0}^{\top}s(z)\big{missing})\prod_{i}\exp\big(\lambda_{i}^{\top}s(z)\big{missing})\approx p^{*}(z).$ (25)
The $i$-th site parameter $\lambda_{i}$ can loosely be interpreted as the natural parameters of a $\mathcal{F}$-approximation to the $i$-th target factor, $\exp\bm{(}\ell_{i}(z)\bm{)}$. EP is also often used to approximate the normalising constant of $p^{*}(z)$ – we have omitted details here, but see e.g. Bishop [3].
In order to update $\lambda_{i}$, EP performs a KL-divergence minimisation between $p(z)$ and a local approximation to the target distribution, consisting of just one “real” factor and $(m-1)$ approximate ones. Specifically,
$\lambda_{i}$ $\leftarrow\operatorname*{arg\,min}_{\lambda_{i}}\mathrm{KL}\left[\bar{p}_{i}(z)\;\middle\|\;p(z;\lambda_{i})\right],$ (26)
where
$\bar{p}_{i}(z)\propto\exp\big(\ell_{i}(z)\big{missing})\prod_{j\neq i}\exp\big(\lambda_{j}^{\top}s(z)\big{missing}),$ (27)
and we have used the notation $p(z;\lambda_{i})$ to make the dependence on $\lambda_{i}$ explicit. It can be easily shown that the solution to (26) is found by moment matching. That is, the KL-divergence is minimised when
$\mathbb{E}_{p(z;\lambda_{i})}[s(z)]$ $=\mathbb{E}_{\bar{p}_{i}(z)}[s(z)].$ (28)
For a minimal exponential family $\mathcal{F}$, there is a unique member of $\mathcal{F}$ with mean parameters $\mathbb{E}_{\bar{p}_{i}(z)}[s(z)]$, and its natural parameters are given by $\nabla A^{*}\bm{(}\mathbb{E}_{\bar{p}_{i}(z)}[s(z)]\bm{)}$. By equating this with the parameters of $p(z)$, we have
$\lambda_{i}$ $=\nabla A^{*}\bm{(}\mathbb{E}_{\bar{p}_{i}(z)}[s(z)]\bm{)}-\eta_{0}-\sum_{j\neq i}\lambda_{j}.$ (29)
Update (29) is the “standard” update of EP, and can be performed either serially or in parallel (over $i$). Often a damping factor $\alpha$ is used, giving
$\lambda_{i}$ $=(1-\alpha)\lambda_{i}+\alpha\left(\nabla A^{*}\bm{(}\mathbb{E}_{\bar{p}_{i}(z)}[s(z)]\bm{)}-\eta_{0}-\sum_{j\neq i}\lambda_{j}\right).$ (30)
In Section 2 we presented EP and several variants as performing some number of “inner” updates, to decrease $L$ with respect to $\lambda_{i}$, with an “outer” update with respect to $\theta$. These updates are given by
Inner update: $\lambda_{i}\leftarrow\lambda_{i}-\alpha\Big{(}\eta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)}\Big{)},$ (31)
Outer update: $\theta\leftarrow\eta_{0}+\sum_{j}\lambda_{j}.$ (32)
Immediately after an outer update we have that $\theta=\eta_{0}+\sum_{j}\lambda_{j}$, and so for $\beta_{i}=1$ we have
$p_{i}(z)$ $\propto\exp\left((\theta-\lambda_{i})^{\top}s(z)+\ell_{i}(z)\right)$
$=\exp\left((\eta_{0}+\sum_{j\neq i}\lambda_{j})^{\top}s(z)+\ell_{i}(z)\right)$
$\propto\bar{p}_{i}(z).$ (33)
We stated in Section 2 that (non-double-loop) EP can be viewed as performing a single inner update per outer update, and so by rolling (31) and (32) into a single update for $\lambda_{i}$, we have
$\lambda_{i}$ $\leftarrow\lambda_{i}-\alpha\Big{(}\eta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{\bar{p}_{i}(z)}[s(z)]\bm{)}\Big{)}$
$=(1-\alpha)\lambda_{i}+\alpha\left(\nabla A^{*}\bm{(}\mathbb{E}_{\bar{p}_{i}(z)}[s(z)]\bm{)}-\eta_{0}-\sum_{j\neq i}\lambda_{j}\right),$ (34)
which is identical to (30).
Appendix C EP inner update derivation
@@
$\nabla_{\lambda_{i}}L(\theta,\lambda_{1},\ldots,\lambda_{m})$ $=\nabla A\left(\eta_{0}+\sum_{j}\lambda_{j}\right)+\beta_{i}\nabla_{\lambda_{i}}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}.$ (35)
@@
$\nabla A\left(\eta_{0}+\lambda_{i}^{\prime}+\sum_{j\neq i}\lambda_{j}\right)$ $=-\beta_{i}\nabla_{\lambda_{i}}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}$
@@
Appendix D Natural gradient view of EP
Let $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ be the member of $\mathcal{F}$ with natural parameter $\smash{\cramped{\eta_{i}}^{(t)}(\lambda_{i})=\eta_{0}+\cramped{\sum}_{j}\cramped{\lambda_{j}}^{(t)}+\alpha^{-1}(\lambda_{i}-\cramped{\lambda_{i}}^{(t)})}$. Proposition 1 states that that the moment-matching updates of EP can be viewed as performing NGD in $L$ with respect to the mean parameters, $\mu_{i}$, of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$. We restate the proposition below, and give a proof. Note that the right hand side of (7) maps the EP update (4) onto mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$. The left hand side of (7) is simply a NGD update in $\mu_{i}$.